# Reliable and efficient inverse analysis using physics-informed neural networks with distance functions and adaptive weight tuning

### v1

@@ 3.3 Inverse problem: incompressible flow around an obstacle
In the computational fluid dynamics community, it is common practice to employ dense and refined meshes near obstacles with complex geometries so that one can accurately capture steep gradients of the flow. Motivated by this, we first compared the PINN’s performance using collocation points taken from the finite element nodes (increased density near the obstacle) and those uniformly drawn from the domain (see Figure 13 (a)). For this preliminary test, we used hard imposition approach without dynamic normalization to focus on the effect of the collocation point distribution. Velocity reconstruction results are presented in Figure 14 (b). Although both are not quite satisfactory, the uniform distribution of collocation points yields better results. This result is fairly natural from Monte Carlo perspective, as each collocation point holds a representative volume to discretize the integrals in the loss function. Indeed, low-discrepancy sequences (e.g., lattice) are known to generally facilitate quicker convergence [62]. For non-uniform distributions, techniques such as volumetric averaging [104] can be incorporated to adjust the representative volume based on collocation point density. However, a comprehensive study of non-uniform collocation points is beyond the scope of this work. Based on this observation, we proceed with uniformly distributed collocation points for the problem with the heart-shaped obstacle.
@@
Figure 14 shows the velocity field, pressure, and estimated kinematic viscosity obtained using the soft and hard impositions with dynamic normalization. From Figure 14 (a), one can observe that the solution obtained with soft imposition, even with dynamic normalization, does not exhibit the shedding structure, resembling a highly viscous or steady flow. This observation aligns with [105], and is attributed to improper boundary condition treatment. Results with the hard imposition clearly reproduce the vortex shedding, and the pressure distribution is also accurately inferred. The kinematic viscosity estimation is illustrated in Figure 14 (b) and summarized in Table 4. One can further confirm that the unobserved viscosity is accurately identified by the combination of hard imposition and dynamic normalization. Notably, neither distance function-based enforcement nor dynamic normalization alone is sufficient; both techniques are necessary to achieve accurate results. Once again, the distance field constructed upon R-functions [50, 51, 52] enables the detection of both outer and interior boundaries, and further extends to non-convex geometries, such as the heart-shaped domain in this problem. This capability is a unique feature of R-function-based distance fields, which approaches in [2, 39, 45, 34, 66] cannot handle, as they implicitly assume convex hull boundaries. Furthermore, Figure 15 illustrates the evolution of the adaptive weights built by bias-corrected dynamic normalization. The weights exhibit stable behavior when combined with the distance functions, which is similar to the behavior of adaptive activation functions observed in the channel flow problems by Sun et al. [39]. These results suggest that hard boundary condition enforcement contributes to the stability of the adaptive weights.
@@ 4 Conclusion
In this study, we investigated the impact of boundary condition imposition techniques on the performance of physics-informed neural networks for both forward and inverse problems of partial differential equations. Although penalty-based soft enforcement is prevalent in the literature, its performance is highly sensitive to the weight parameter, requiring very careful tuning [10, 40]. In contrast, distance function-based hard enforcement directly encodes the boundary conditions into the approximate function structure, ensuring that the solution inherently satisfies the prescribed conditions. In addition, we have reviewed the properties of R-functions [50, 51, 52, 64] while exploring the use of distance functions in the context of boundary value problems. In the framework of R-functions, the distance field is normalized, providing a smooth and differentiable representation of distance and unit normal vectors on the boundaries. These properties are crucial for applications to homogeneous/inhomogeneous Dirichlet, Neumann, and mixed boundary conditions [55, 56, 57, 46, 58] and enable the detection of arbitrary polygonal and non-convex boundaries, overcoming limitations of other (non-normalized) distance function-based approaches [2, 39, 45, 66].
@@
The effectiveness of distance-based hard imposition was first verified with a forward problem of the Poisson equation and subsequently extended to inverse problems of incompressible flow in a cavity and around an obstacle. For forward problems, soft imposition approach quickly reached a plateau, which is consistent result with [39]. In contrast, hard imposition outperformed soft enforcement across various activation functions, achieving faster convergence and lower error. The same architecture was then applied to inverse problems, however, we found that the sole application of hard imposition to explicitly enforce the boundary conditions was insufficient to achieve accurate inverse analysis. To address this, we introduced dynamic normalization [47], which adaptively adjusts the weights assigned to the data and PDE losses, and combined it with hard imposition. The proposed framework, incorporating hard imposition (based on distance functions) and dynamic normalization (adaptive weight tuning), provided a robust and efficient inverse analysis approach. This study underscores the importance of boundary condition treatment and the efficacy of adaptive weight tuning in improving the performance of PINNs, particularly for inverse problems.
The presented method, however, is not without limitations. For example, the neural network structure is a simple MLP, and advanced architectures and preprocessing techniques, such as basis transformation [33, 34], domain decomposition [35, 37, 36], and mixed-variable formulations [106, 84] could be explored to capture the complex flow structures more effectively.
@@ Acknowledgements
This work was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP23KK0182, JP23K17807, JP23K26356, JP23K24857, JP23KJ1685, and Japan Science and Technology Agency (JST) Support for Pioneering Research Initiated by the Next Generation (SPRING) Grant Number JPMJSP2136.
@@ Appendix A Numerical study on several improvement techniques
There have been many studies on adaptive activation functions in the literature of neural networks [107, 86]. Similarly, Jagtap et al. [83, 88] proposed adaptive activation functions in the context of PINNs, in which the forward pass is modified as follows:
@@
where a learnable scaling factor $s^{(l)}$ is introduced in each layer. Equation (A.1) is known as layer-wise locally adaptive activation function (L-LAAF) [88]. With this modification, the set of trainable parameters becomes $\bm{\theta}=\left\{\mathbf{W}^{(l)},\bm{b}^{(l)},s^{(l)}\right\}_{l=1}^{L}\setminus\left\{s^{(L)}\right\}$. The core idea behind this modification is to adaptively adjust the slope of the activation functions. Specifically, backpropagated gradients are amplified if $s^{(l)}\geq 1$ and attenuated if $s^{(l)}<1$. To encourage $s^{(l)}$ to increase, thereby accelerating optimization, Jagtap et al. [88] introduced the following slope recovery term into the loss function:
@@
Yu et al. [25] proposed a gradient enhancement technique, termed gPINN, where an additional loss term is introduced to penalize the derivatives of the PDE residual so that they approach zero. In particular, they introduced the following gradient enhancement term:
[…]
While higher-order derivatives can be included to further enforce smoothness, the computational cost increases rapidly with the order of differentiation. Moreover, the effectiveness of gradient enhancement is highly sensitive to the weight $\lambda_{\text{GE}}$, and improper tuning of $\lambda_{\text{GE}}$ may lead to degraded performance, whereas gPINN could be effective for scarce data and help prevent oscillation due to its smoothing effect [108, 109].
@@
McClenny & Braga-Neto [90] proposed self-adaptive PINN (SA-PINN), in which a set of learnable weights is introduced to adaptively determine the importance of each collocation point. In SA-PINN, the loss function is modified as follows:
@@
where $\bm{\lambda}=\left\{\bm{\lambda}_{\text{PDE}},\bm{\lambda}_{\text{DBC}},\bm{\lambda}_{\text{NBC}},\bm{\lambda}_{\text{Data}}\right\}$ is a set of non-negative learnable weights assigned to each collocation point. $m:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a smooth, strictly monotonically increasing function, typically chosen as $m\left(\cdot\right)=\max\left(0,\cdot\right)^{2}$ [90, 110]. The training of the self-adaptive PINN is formulated as $\min_{\bm{\theta}}\max_{\bm{\lambda}}\mathcal{L}\left(\bm{\theta},\bm{\lambda}\right)$, where $\bm{\theta}$ is updated using gradient descent, while $\bm{\lambda}$ is updated by gradient ascent algorithm:
@@
Here, $\eta_{\bm{\theta}}$ and $\eta_{\bm{\lambda}}$ are the learning rates for $\bm{\theta}$ and $\bm{\lambda}$, respectively. In [90], authors recommend using a larger learning rate for $\bm{\lambda}$ compared to $\bm{\theta}$, i.e., $\eta_{\bm{\lambda}}=c_{\text{SA}}\eta_{\bm{\theta}}$, for some hyperparameter $c_{\text{SA}}\geq 1$. Self-adaptive PINN has been applied to various domains, such as infection dynamics [110] and seismic wave modeling [111].
Table A.1: Poisson equation with mixed boundary conditions: homogeneous Neumann condition case (described in Section 3.1). Comparison of L-LAAF [88], gPINN [25], and SA-PINN [90]. Relative $L^{2}$ error (mean and standard error) using GELU activation function. All values are scaled by $10^{-3}$ for the consistency with Table 1.
@@
Table A.1 summarizes the relative $L^{2}$ error with L-LAAF [88], gPINN [25], and SA-PINN [90] for the Poisson equation problem with inhomogeneous Dirichlet and homogeneous Neumann conditions. The problem setup is the same as that in Section 3.1, and the hyperparameters are selected based on the numerical experiments in the original papers and their recommendations described therein. For SA-PINN, we considered two distinct strategies: one where both $\bm{\theta}$ and $\bm{\lambda}$ are updated throughout the optimization (SA-PINN), and the other where a two-stage optimization is employed (SA-PINN). In the latter case, $\bm{\theta}$ and $\bm{\lambda}$ were jointly optimized for the first half of the total training epochs, after which $\bm{\lambda}$ was frozen, and only $\bm{\theta}$ was updated for the remaining epochs, following [90]. The results suggest that these methods do not consistently yield substantial improvements, and their effectiveness is highly sensitive to the hyperparameters.
@@ Appendix B Numerical study on normalization orders in distance functions
In [46, 58], the behavior of the Laplacian of the ADFs is discussed analytically. Here, we numerically investigate the influence of the normalization order $m$ of the distance functions on the convergence of PINN solutions. Specifically, we consider the Poisson equation problem (described in Section 3.1) as an example. Figure B.1 illustrates the error with different activation functions and normalization orders. The results, summarized in Table B.1, indicate a clear trend: as the normalization order $m$ increases, the convergence of PINN solutions deteriorates for all tested activation functions, although the extent of the degradation varies.
@@
$\nabla^{2}\tilde{u}$ $=\nabla\cdot\nabla\left(\phi^{(m)}\hat{u}\right)$ (B.1)
$=\nabla\cdot\left(\hat{u}\nabla\phi^{(m)}+\phi^{(m)}\nabla\hat{u}\right)$ (B.2)
@@
[…]

### v3

@@ 3.3 Inverse problem: incompressible flow around an obstacle
In the context of computational fluid dynamics, it is common practice to employ dense and refined meshes near obstacles with complex geometries so that one can accurately capture steep gradients of the flow. Motivated by this, we first compared the PINN’s performance using collocation points taken from the finite element nodes (increased density near the obstacle) and those uniformly drawn from the domain (see Figure 14 (a)). For this preliminary test, we used hard imposition approach without dynamic normalization to focus on the effect of the collocation point distribution. Velocity reconstruction results are presented in Figure 14 (b). Although both are not quite satisfactory, the uniform distribution of collocation points yields better results. This result is fairly natural from Monte Carlo perspective, as each collocation point holds a representative volume to discretize the integrals in the loss function. Indeed, low-discrepancy sequences (e.g., lattice) are known to generally facilitate quicker convergence [Matsubara2023GLT]. For non-uniform distributions, techniques such as volumetric averaging [Song2024VWPINN] can be incorporated to adjust the representative volume based on collocation point density. However, a comprehensive study of non-uniform collocation points is beyond the scope of this work. Based on this observation, we proceed with uniformly distributed collocation points for the problem with the heart-shaped obstacle.
@@
Figure 15 shows the velocity field, pressure, and estimated kinematic viscosity obtained using the soft and hard impositions with dynamic normalization. From Figure 15 (a), one can observe that the solution obtained with soft imposition, even with dynamic normalization, does not exhibit the shedding structure, resembling a highly viscous or steady flow. This observation aligns with [Chuang2022ExperienceReport], and is attributed to improper boundary condition treatment. Results with the hard imposition clearly reproduce the vortex shedding, and the unobserved pressure is also accurately inferred. The kinematic viscosity estimation is illustrated in Figure 15 (b) and summarized in Table 5. One can further confirm that the unobserved viscosity is accurately identified by the integration of hard imposition and dynamic normalization. Notably, neither distance function-based enforcement nor dynamic normalization alone is sufficient; both techniques are necessary to achieve accurate results (see Table 5 and Figure 14 (b) (right)). Once again, the distance field constructed upon R-functions [Sheiko1982, Rvachev2001, Shapiro2007] enables the detection of both outer and interior boundaries, and further extends to non-convex geometries, such as the heart-shaped domain in this problem. This capability is a unique feature of R-function-based distance fields, whereas approaches in [Lagaris1998, Sun2020Surrogate, Lu2021HardDesign, Tang2023PolyPINN] cannot handle, as they implicitly assume convex hull boundaries. Furthermore, Figure 16 illustrates the evolution of the adaptive weights built by bias-corrected dynamic normalization. The weights exhibit stable behavior when combined with the distance functions, which is similar to the behavior of adaptive activation functions observed in the channel flow problems by Sun et al. [Sun2020Surrogate]. These results suggest that hard boundary condition enforcement contributes to the stability of the adaptive weights.
@@ 4 Conclusion
In this study, we investigated the impact of boundary condition imposition techniques on the performance of physics-informed neural networks for both forward and inverse problems of partial differential equations. Although penalty-based soft enforcement is prevalent in the literature, its performance is highly sensitive to the weight parameter, requiring very careful tuning [E2018DRM, Chen2020DGMvsDRM]. In contrast, distance function-based hard enforcement directly encodes the boundary conditions into the approximate function structure, ensuring that the solution inherently satisfies the prescribed conditions. In addition, we have reviewed the properties of exact and approximate distance functions [Biswas2004] while exploring the use of distance functions in the context of boundary value problems. In the framework of R-functions, the distance field can be normalized, providing a smooth and differentiable representation of distance and unit normal vectors on the boundaries. These properties are crucial for applications to homogeneous/inhomogeneous Dirichlet, Neumann, and mixed boundary conditions [Shapiro2002SAGE, Tsukanov2003MeshfreeFluid, Millan2015, Sukumar2022MixedBC, Berrone2023BCEnforcement] and enable the detection of arbitrary polygonal and non-convex boundaries, overcoming limitations of other (naïve and non-normalized) distance function-based approaches such as those presented in [Lagaris1998, Lu2021HardDesign, Li2024HardAdvcDiff].
@@
The effectiveness of normalized distance-based hard imposition was first verified with a forward problem of the Poisson equation and subsequently extended to inverse problems of incompressible flow in a cavity and around an obstacle. For forward problems, soft imposition approach quickly reached a plateau, which is consistent result with [Sun2020Surrogate]. In contrast, hard imposition outperformed soft enforcement across various activation functions, achieving faster convergence and lower error. The same architecture was then applied to inverse problems, however, we found that the sole application of hard imposition to explicitly enforce the boundary conditions was insufficient to achieve accurate inverse analysis. To address this, we introduced dynamic normalization [Deguchi2023DynNorm], which adaptively adjusts the weights assigned to the data and divergence losses, and combined it with hard imposition. The proposed framework, incorporating hard imposition (R-function-based normalized distance) and dynamic normalization (bias-corrected adaptive weight), provided a robust and efficient inverse analysis approach. This study underscores the importance of boundary condition treatment and the efficacy of adaptive weight tuning in improving the performance of PINNs, particularly for inverse problems.
In this study, we have demonstrated the applicability of the proposed method to steady and unsteady problems. The methodology can be further extended to initial-boundary value problems by interpreting the initial condition as a special case of Dirichlet condition on the spatio-temporal plane [Sun2020Surrogate, Berrone2023BCEnforcement, Kharazmi2021hpVPINN, Matsubara2023GLT] and extending the R-function-based distance fields to spatio-temporal domains (e.g., [Sanchez2015SpaceTimeRfunction, Mulroy2022RfunctionSoftRobots]). Notably, for time-dependent problems where the solution exhibits dynamic, multi-scale, or turbulent behavior, one could consider loss function formulations that explicitly account for temporal causality [Wang2024Causality] to further enhance the network to learn the temporal evolution.
The presented method can be further improved by exploring various aspects. For instance, the neural network structure employed was a simple MLP. Future investigations could explore the integration with advanced network architectures and preprocessing techniques, such as basis transformation [Wang2021FourierFeature, Tang2023PolyPINN, Wang2024PirateNets], domain decomposition [Jagtap2020XPINN, Moseley2023FBPINN, Li2023D3M], and mixed-variable formulations [Rao2020TAML, Haghighat2021SolidMechanics], to more effectively capture complex flow structures. Furthermore, future work could explore the integration of different architectures, such as convolutional neural networks to efficiently capture spatial features [Gao2021PhyGeoNet] or attention mechanisms [Vaswani2017Transformer] to automatically detect important regions of the domain [McClenny2023SA]. The applicability of the current R-function-based hard imposition primarily focused on fixed boundaries. Extending this framework to handle domain changes, such as moving boundaries, presents another avenue for future research, which could potentially be addressed by constructing time-varying distance functions [Sanchez2015SpaceTimeRfunction].
@@ Acknowledgements
This work was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP25KJ0026, JP23KK0182, JP23K17807, JP23K26356, JP23K24857, JP23KJ1685, and Japan Science and Technology Agency (JST) Support for Pioneering Research Initiated by the Next Generation (SPRING) Grant Number JPMJSP2136.
@@ Appendix A Numerical study on several improvement techniques
There have been many studies on adaptive activation functions in the literature of neural networks [Chen1996AGtanh, Ramachandran2017Swish]. Similarly, Jagtap et al. [Jagtap2020GAAF, Jagtap2020LAAF] proposed adaptive activation functions in the context of PINNs, in which the forward pass is modified as follows:
@@
where a learnable scaling factor $s^{(l)}$ is introduced in each layer. Equation (A.1) is known as layer-wise locally adaptive activation function (L-LAAF) [Jagtap2020LAAF]. With this modification, the set of trainable parameters becomes $\bm{\theta}=\left\{\mathbf{W}^{(l)},\bm{b}^{(l)},s^{(l)}\right\}_{l=1}^{L}\setminus\left\{s^{(L)}\right\}$. The core idea behind this modification is to adaptively adjust the slope of the activation functions. Specifically, backpropagated gradients are amplified if $s^{(l)}\geq 1$ and attenuated if $s^{(l)}<1$. To encourage $s^{(l)}$ to increase, thereby accelerating optimization, Jagtap et al. [Jagtap2020LAAF] introduced the following slope recovery term into the loss function:
@@
Yu et al. [Yu2022GradEnhc] proposed a gradient enhancement technique, termed gPINN, where an additional loss term is introduced to penalize the derivatives of the PDE residual so that they approach zero. In particular, they introduced the following gradient enhancement term:
$\mathcal{L}_{\text{GE}}=\sum_{k=1}^{d}\int_{\Omega}\left|\frac{\partial}{\partial x_{k}}\left(-\nabla^{2}\hat{u}\left(\bm{x};\bm{\theta}\right)-f\left(\bm{x}\right)\right)\right|^{2}\,d\bm{x}.$ (A.3)
While higher-order derivatives can be included to further enforce smoothness, the computational cost increases rapidly with the order of differentiation. Moreover, the effectiveness of gradient enhancement is highly sensitive to the weight $\lambda_{\text{GE}}$, and improper tuning of $\lambda_{\text{GE}}$ may lead to degraded performance, whereas gPINN could be effective for scarce data and help prevent oscillation due to its smoothing effect [Mohammadian2023gPINN, Eshkofti2023gPINN].
@@
McClenny & Braga-Neto [McClenny2023SA] proposed self-adaptive PINN (SA-PINN), in which a set of learnable weights is introduced to adaptively determine the importance of each collocation point. In SA-PINN, the loss function is modified as follows:
@@
where $\bm{\lambda}=\left\{\bm{\lambda}_{\text{PDE}},\bm{\lambda}_{\text{DBC}},\bm{\lambda}_{\text{NBC}},\bm{\lambda}_{\text{Data}}\right\}$ is a set of non-negative learnable weights assigned to each collocation point. $m:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a smooth, strictly monotonically increasing function, typically chosen as $m\left(\cdot\right)=\max\left(0,\cdot\right)^{2}$ [McClenny2023SA, Linka2022BPINNCovid]. The training of the self-adaptive PINN is formulated as $\min_{\bm{\theta}}\max_{\bm{\lambda}}\mathcal{L}\left(\bm{\theta},\bm{\lambda}\right)$, where $\bm{\theta}$ is updated using gradient descent, while $\bm{\lambda}$ is updated by gradient ascent algorithm:
@@
Here, $\eta_{\bm{\theta}}$ and $\eta_{\bm{\lambda}}$ are the learning rates for $\bm{\theta}$ and $\bm{\lambda}$, respectively. In [McClenny2023SA], authors recommend using a larger learning rate for $\bm{\lambda}$ compared to $\bm{\theta}$, i.e., $\eta_{\bm{\lambda}}=c_{\text{SA}}\eta_{\bm{\theta}}$, for some hyperparameter $c_{\text{SA}}\geq 1$. Self-adaptive PINN has been applied to various domains, such as infection dynamics [Linka2022BPINNCovid] and seismic wave modeling [Ding2023SAPINN_Seamic].
(a) L-LAAF
(b) gPINN
(c) SA-PINN
(d) SA-PINN
Figure A.1: Poisson equation with mixed boundary conditions: homogeneous Neumann condition case (described in Section 3.1). Relative $L^{2}$ error of the approximate solution with L-LAAF [Jagtap2020LAAF], gPINN [Yu2022GradEnhc], and SA-PINN [McClenny2023SA] using GELU activation function. The dashed lines represent the mean, and the shaded areas represent the standard error over 5 independent runs with i.i.d. initialization. Notably, even with these improvement techniques, soft imposition quickly faces a plateau, as observed in Figure 7 (a).
Table A.1: Poisson equation with mixed boundary conditions: homogeneous Neumann condition case (described in Section 3.1). Comparison of L-LAAF [Jagtap2020LAAF], gPINN [Yu2022GradEnhc], and SA-PINN [McClenny2023SA]. Relative $L^{2}$ error (mean and standard error) using GELU activation function. All values are scaled by $10^{-3}$ for the consistency with Table 2.
@@
Figure A.1 and Table A.1 summarizes the relative $L^{2}$ error with L-LAAF [Jagtap2020LAAF], gPINN [Yu2022GradEnhc], and SA-PINN [McClenny2023SA] for the Poisson equation problem with inhomogeneous Dirichlet and homogeneous Neumann conditions. The problem setup is the same as that in Section 3.1, and the hyperparameters are selected based on the numerical experiments in the original papers and their recommendations described therein. For SA-PINN, we considered two distinct strategies: one where both $\bm{\theta}$ and $\bm{\lambda}$ are updated throughout the optimization (SA-PINN), and the other where a two-stage optimization is employed (SA-PINN). In the latter case, $\bm{\theta}$ and $\bm{\lambda}$ were jointly optimized for the first half of the total training epochs, after which $\bm{\lambda}$ was frozen, and only $\bm{\theta}$ was updated for the remaining epochs, following [McClenny2023SA]. The results suggest that these methods do not consistently yield substantial improvements, and their effectiveness is highly sensitive to the hyperparameters.
@@ Appendix B Numerical study on normalization orders in distance functions
In [Sukumar2022MixedBC, Berrone2023BCEnforcement], the behavior of the Laplacian of the ADFs is discussed analytically. Here, we numerically investigate the influence of the normalization order $m$ of the distance functions on the convergence of PINN solutions. Specifically, we consider the Poisson equation problem (described in Section 3.1) as an example. Figure B.1 illustrates the error with different activation functions and normalization orders. The results, summarized in Table B.1, indicate a clear trend: as the normalization order $m$ increases, the convergence of PINN solutions deteriorates for all tested activation functions, although the extent of the degradation varies.
@@
[…]
@@
where $\left(\cdot,\cdot\right)$ denotes the inner product. The presence of $\nabla^{2}\phi^{(m)}$ in $\nabla^{2}\tilde{u}$ suggests a high possibility that $\nabla^{2}\tilde{u}$ inherits the oscillatory nature of $\nabla^{2}\phi^{(m)}$, potentially resulting in the degraded convergence observed in Figure B.1 and Table B.1. This implies that the precise representation of the distance (i.e., higher $m$) is not necessarily advantageous or beneficial; rather, the normalization order $m$ should be kept moderate so as not to require $\hat{u}$ to compensate for the near-boundary oscillations $\nabla^{2}\phi^{(m)}$ may induce, especially when the second derivative appears in the problem of interest.